# Understanding Software Engineering Agents: A Study of Thought-Action-Result Trajectories

### v1

@@ Abstract
Large Language Model (LLM)-based agents are increasingly employed to automate complex software engineering tasks such as program repair and issue resolution. These agents operate by autonomously generating natural language thoughts, invoking external tools, and iteratively refining their solutions. Despite their widespread adoption, the internal decision-making processes of these agents remain largely unexplored, limiting our understanding of their operational dynamics and failure modes. In this paper, we present a large-scale empirical study of the thought-action-result trajectories of three state-of-the-art LLM-based agents: RepairAgent, AutoCodeRover, and OpenHands. We unify their interaction logs into a common format, capturing 120 trajectories and 2822 LLM interactions focused on program repair and issue resolution. Our study combines quantitative analyses of structural properties, action patterns, and token usage with qualitative assessments of reasoning coherence and feedback integration. We identify key trajectory characteristics such as iteration counts and token consumption, recurring action sequences, and the semantic coherence linking thoughts, actions, and their results. Our findings reveal behavioral motifs and anti-patterns that distinguish successful from failed executions, providing actionable insights for improving agent design, including prompting strategies, failure diagnosis, and anti-pattern detection. We release our dataset and annotation framework to support further research on transparent and robust autonomous software engineering agents.
@@ I Introduction
Beyond querying LLMs with fixed prompt templates and within hard-coded algorithms, LLM agents have recently emerged as another promising approach to software engineering automation. Such agents are autonomous systems that use LLMs to iteratively reason about a problem, invoke existing tools, and adapt based on outputs produced by these tools [13]. For example, LLM agents can autonomously fix bugs and other issues [14, 15, 16, 17], write tests to reproduce issues reported in natural language [18, 19, 20], and automate complex project setups [21, 22]. To address a given task, an LLM agent autonomously plans and executes a sequence of actions, leveraging its reasoning capabilities and external tools. Agents in software engineering typically use tools such as code editors, compilers, debuggers, and test runners to interact with the code base and the environment.
While the success of LLM agents in software engineering tasks is promising, the underlying decision-making processes remain largely opaque. However, as their adoption increases, a critical question emerges: How do LLM agents reach their solutions, and why do they (not) work well on specific tasks? Answering this question is crucial for understanding the strengths and limitations of LLM agents, as well as for improving their performance and reliability. In particular, insights into the usage of tools could help to improve the design of LLM agents and their integration with existing software engineering tools. Likewise, insights into the reasoning processes of agents could reveal common pitfalls and best practices, which will provide a basis for improving future LLM agents.
Understanding how LLM agents address software engineering tasks is important but also challenging. On the upside, LLM agents provide detailed logs that document their reasoning and the actions they take. We call these logs agent trajectories, which we define as a sequence of thought-action-result triples, where each triple consists of a reasoning step, or thought, expressed by an LLM, an external action (e.g., invoking a compiler), and the result of that action (e.g., the compiler output). While the availability of agent trajectories is a promising starting point, there currently is no established methodology to systematically study them.
@@
  • RQ1: What are the overall properties of trajectories, such as the number of iterations and consumed LLM-tokens?
  • RQ2: What kinds of actions do agents perform and what sequential patterns emerge within trajectories?
@@
To address these questions, we introduce a novel methodology that analyzes LLM-based software engineering agents through their thought-action-result trajectories. We begin by collecting logs from the three studied agents (RepairAgent, AutoCodeRover, OpenHands) and convert them into a unified trajectory format. We then compute statistics to examine trajectory properties and how to relate to success or failure of an agent (RQ1). Next, we manually categorize agent actions, and mine frequent action sequences (RQ2). Finally, to study reasoning coherence (RQ3), we apply open coding to label semantic relationships between thoughts, actions, and results. This mixed-method approach enables us to extract both quantitative insights and qualitative interpretations of agent behavior. Our study leads to several key findings:
@@
  • Successful action sequences balance exploration, explanation, fix generation, and testing; failures exhibit repetitive, non-adaptive cycles and debugging anti-patterns like repeated identical actions without follow-up or fix generation without intermediate testing.
  • Semantic alignment between thoughts and actions is essential; even rare misalignments strongly correlate with failure or increased computational cost, underscoring the need for explicit verification mechanisms (e.g., self-reflection or critique) to ensure coherence.
@@
    Methodology. We introduce a novel trajectory-based modeling and analysis capturing the interplay of thoughts, actions, and results in SE agents.
@@
    Empirical results. We provide the first systematic study of agent trajectories across multiple state-of-the-art SE agents, revealing characteristic patterns and challenges.
@@ II-A Data Collection
We collect trajectory logs from three SOTA SE agents:
  • RepairAgent [^1] https://github.com/sola-st/RepairAgent [14]: An autonomous agent for program repair evaluated on 835 bugs of Defects4J [32].
@@
To ensure a diverse yet manageable dataset for detailed manual annotation, we randomly sample 40 trajectories per agent. Given the average success rate of approximately 27% across the three agents, we proportionally sample about 10 successful trajectories per agent. Each trajectory captures a complete bug repair or issue resolution through iterative agent-environment interactions. Our dataset thus comprises 120 trajectories totaling 2822 iterations across all agents.
@@ II-C Statistical Analysis
The trajectory length reflects task complexity, with easier tasks completed in fewer iterations. To manage computational effort, agents limit trajectory length, e.g., RepairAgent caps at 40 iterations, while OpenHands allows up to 100.
@@
[…]
@@
We define success as in reported in the original evaluation of the agents, i.e., whether the agent produced a valid patch or solution to the task. Specifically, for AutoCodeRover and OpenHands, we consider a trajectory successful if the agent produces a patch that passes the tests executed by SWE-bench [33]. For RepairAgent, we consider a trajectory successful if the produced patch passes the tests executed by Defects4J [32] (called “plausible”) and manual inspection confirms that the patch is equal or semantically to the developer-produced patch (called “correct”).
@@ II-D Categorizing Actions
To categorize the actions, we first map known tools used by agents such as search tools into the respective category (e.g, Search). Then, we manually inspect the remaining actions that we were unable to map automatically (e.g, a terminal command). The labeling is initially performed by one author, followed by a discussion of ambiguous and uncertain cases to resolve them collaboratively among multiple authors. A small subset of actions (8.3%) does not fit into any of the above categories, such as internal state transitions in RepairAgent and LLM-suggested actions that AutoCodeRover fails to parse. We exclude these cases from subsequent analyses.
@@ II-E Mining Sequential Action Patterns
We analyze sequences of categorized actions in agent trajectories to identify recurrent decision-making patterns. Each trajectory is represented as a sequence of action labels, e.g., Explore $\rightarrow$ Locate $\rightarrow$ Generate Fix $\rightarrow$ Reproduce $\rightarrow$ Run Tests. We study these sequences by mining for frequently occurring subsequences. First, we implement a lightweight Python script to extract fixed‑length $n$‑grams and compute their frequencies, choosing $n$ large enough to avoid trivial short sequences but small enough to retain sufficient diversity and easy-to-analyze patterns; from our experiments, $n\in\{4,5,6\}$ performs best, and for this study we set $n=4$, which yields the richest set of recurring patterns. Next, for each agent we identify the top‑frequency 4‑grams and analyze how their relative frequencies differ between successful and failed tasks. Finally, we flag patterns overrepresented in failing trajectories as well as rare or unusual 4‑grams that may indicate nonstandard or suboptimal workflows.
@@ II-F Labeling Semantic Relationships
Thought $t_{i}$rightarrow action $t_{i}$ (within one iteration):
@@
Thought $t_{i}$rightarrow thought $t_{i+1}$ (across consecutive iterations):
@@
Action $a_{i}$rightarrow action $a_{i+1}$ (across consecutive iterations):
@@
Result $r_{i}$rightarrow thought $t_{i+1}$ (across consecutive iterations):
@@
Result $r_{i}$rightarrow action $a_{i+1}$ (across consecutive iterations):
@@
These relationships capture how the agent’s results influence its subsequent actions. For example, we label a pair of result and action as Informative when the result provides information that is incorporated into the next action (e.g, api name found through search).
@@
The relationships in Table I are the result of labeling 40 trajectories from each of the three agents, which totals to 14K pairs of trajectory components. Initially, some of the authors label each pair, while discussing ambiguous and uncertain cases with other authors to resolve them collaboratively.
@@ III-A RQ1: Trajectory-Level Properties
We investigate trajectory properties aiming to understand how their iterative processes differ in length and complexity. We answer this by quantitatively analyzing the distribution of iteration counts and tokens consumption.
Figure 2: Comparative trajectory length between successful and unsuccessful trajectories (y-axis differs between the agents).
@@
The figure also compares trajectory lengths between successful and unsuccessful instances, revealing agent-specific patterns. For RepairAgent, unsuccessful cases have longer trajectories (mean 40) than successful ones (mean 22), indicating higher effort on difficult tasks, consistent with its design to continue attempts until a fix is found or early termination is triggered. OpenHands shows a similar trend, though a bit less, with some unsuccessful trajectories reaching the 100-iteration cap. For AutoCodeRover, successful and unsuccessful trajectories have similar average lengths, but unsuccessful ones exhibit higher variability and outliers, often due to internal errors like Response Parse Error. AutoCodeRover terminates after patch application without integrating tests into its workflow.
@@
Figure 3: Input, Output, and Total tokens over successful trajectories and unsuccessful ones (y-axis is different for each agent).
@@
  • AutoCodeRover (ACR): Trajectories are consistently short and efficient; failures often involve Error Response Parsing sequences that modestly increase token usage.
@@ III-B1 Overall Action Usage
Figure 4(a) compares the overall action categories distribution. Across all agents, the most frequent actions are Generate Fix (23%), Run tests (20%), Explain (20%), and Explore (18%). While these core actions are common, each agent exhibits distinct patterns in action selection and sequencing.
RepairAgent tends to emphasize a balanced approach incorporating search and exploration, fix generation, and comprehensive testing. In contrast, OpenHands places particular emphasis on a test-driven methodology, including the generation of issue-reproducing tests, which is distinctive to OpenHands’s workflow. AutoCodeRover operates through a structured search–locate–fix process and typically omitting explicit Reproduce and Test steps in its main pipeline. It is worth noting that a subsequent extension, SpecRover [36], introduces dedicated reproduction and testing phases.
@@ III-B2 Action Usage Over Task Progress
To analyze behavior evolution, we normalize iterations $i\in[1..N]$ of a trajectory to percentage progress. Figures 4(b)4(d) depict the proportion of actions per category at each progress point (e.g., 20%). In RepairAgent (Fig. 4(b)), the initial 0–20% of the trajectory is evenly split between Fix Generation plus Running Tests ( 50%) and the remaining categories (Explore, Explain, Search, Locate). From 20% to 100%, Fix Generation and Running Tests dominate ( 70%). OpenHands (Fig. 4(c)) starts with  50% Explore and Search in the first 20%, then shifts towards increased Fix Generation, Reproduce, and Running Tests, which constitute  75% in the last 20%. AutoCodeRover (Fig. 4(d)), with median $N=5$, spends the first 50% on Search, Locate, and Explain, then transitions to Locate (30%) and Generate Fix (20%) in the latter half. Failing trajectories include many ErrorResponse events, omitted from the plot.
@@ III-B3 Sequence Patterns
To complement the time-distributed action plots, Figures 5, 6, 7 presents a visualization of the top most frequent 4-gram sequences executed by each agent, with each row representing a distinct sequence and its corresponding frequency in successful and unsuccessful trajectories.
In the case of RepairAgent, successful 4-grams prominently feature cycles that alternate between Generate Fix and exploratory actions such as Explain or Explore, with little dominance by any single pattern, as shown in Figure 5. In contrast, unsuccessful trajectories often show a strong concentration in repetitive Generate Fix cycles as shown in the last row of Figure 5 (e.g., multiple consecutive Generate Fix actions).
For OpenHands (Figure 6), the most frequent successful sequences balance between context gathering, patch generating, and testing (generate and run). Unsuccessful trajectories, on the other hand, show more repetitive cycles of exploration and search actions (e.g, first rows in the figure), or sequences of repeated testing and fixing steps without interleaving actions. This highlights the importance balancing between different actions and having strategic shift in debugging phases instead of continuous repetition of same actions.
@@
Figure 5: Top 11 sequences of actions 4-grams (RepairAgent).
Figure 6: Top 12 sequences of actions 4-grams (OpenHands).
Figure 7: Top 11 sequences of actions 4-grams (AutoCodeRover).
@@
1) Repeated identical actions without follow-up: For instance, repeated Search actions occur across all three agents. While some repetitions arise from unsuccessful prior searches, others neglect to act on found results. RQ3 elaborates on this no-influence phenomenon.
2) Repeated “Generate Fix” without testing: Observed mainly in OpenHands and AutoCodeRover, unlike RepairAgent which mandates testing after fix generation. Testing is crucial to identify necessary patch modifications to understand if a fix works and why.

### v2

@@ Abstract
Large Language Model (LLM)-based agents are increasingly employed to automate complex software engineering tasks, such as program repair and issue resolution. These agents operate by autonomously generating natural language thoughts, invoking external tools, and iteratively refining their solutions. Despite their widespread adoption, the internal decision-making processes of these agents remain largely unexplored, limiting our understanding of their operational dynamics and failure modes. In this paper, we present a large-scale empirical study of the thought-action-result trajectories of three state-of-the-art LLM-based agents: RepairAgent, AutoCodeRover, and OpenHands. We unify their interaction logs into a common format, capturing 120 trajectories and 2,822 LLM interactions focused on program repair and issue resolution. Our study combines quantitative analyses of structural properties, action patterns, and token usage with qualitative assessments of reasoning coherence and feedback integration. We identify key trajectory characteristics, such as iteration counts and token consumption, recurring action sequences, and the semantic coherence of thoughts, actions, and their results. Our findings reveal behavioral motifs and anti-patterns that distinguish successful from failed executions, providing actionable insights for improving agent design, including prompting strategies, failure diagnosis, and anti-pattern detection. We release our dataset and annotation framework to support further research on transparent and robust autonomous software engineering agents.
@@ I Introduction
Beyond querying LLMs with fixed prompt templates and within hard-coded algorithms, LLM agents have recently emerged as another promising approach to software engineering automation. Such agents are autonomous systems that use LLMs to iteratively reason about a problem, invoke existing tools, and adapt based on outputs produced by these tools [13]. For example, LLM agents can autonomously fix bugs and other issues [14, 15, 16, 17], write tests to reproduce issues reported in natural language [18, 19, 20], and automate complex project setups [21, 22]. To address a given task, an LLM agent autonomously plans and executes a sequence of actions, leveraging its reasoning capabilities and external tools. Agents in software engineering typically use tools to interact with the code base and the environment, such as code editors, compilers, debuggers, and test runners.
While the success of LLM agents in software engineering is promising, the underlying decision-making processes remain largely opaque. However, as their adoption increases, a critical question emerges: How do LLM agents reach their solutions, and why do they (not) work well on specific tasks? Answering this question is crucial for understanding the strengths and limitations of LLM agents, as well as for improving their performance and reliability. In particular, insights into the usage of tools could help to improve the design of LLM agents and their integration with existing software engineering tools. Likewise, insights into the reasoning processes of agents could reveal common pitfalls and best practices, which will provide a basis for improving future LLM agents.
Understanding how LLM agents address software engineering tasks is important but also challenging. On the upside, LLM agents provide detailed logs that document their reasoning and the actions they take. We call these logs agent trajectories, which we define as a sequence of thought-action-result triples, where each triple consists of a reasoning step, or thought expressed by an LLM, an external action (e.g., invoking a compiler), and the result of that action (e.g., the compiler output). While the availability of agent trajectories is a promising starting point, there currently is no established methodology to systematically study them.
@@
  • RQ1: What are the overall properties of trajectories, such as the number of iterations and consumed LLM tokens?
  • RQ2: What kinds of actions do agents perform, and what sequential patterns emerge within trajectories?
@@
To address these questions, we introduce a novel methodology that analyzes LLM-based software engineering agents through their thought-action-result trajectories. We begin by collecting logs from the three studied agents (RepairAgent, AutoCodeRover, OpenHands) and convert them into a unified trajectory format. We then compute statistics to examine trajectory properties and how they relate to success or failure of an agent (RQ1). Next, we manually categorize agent actions and mine frequent action sequences (RQ2). Finally, to study reasoning coherence (RQ3), we apply open coding to label semantic relationships between thoughts, actions, and results. This mixed-method approach enables us to extract both quantitative insights and qualitative interpretations of agent behavior. Our study leads to several key findings:
@@
  • Successful action sequences balance exploration, explanation, fix generation, and testing. In contrast, failures exhibit repetitive, non-adaptive cycles, such as repeated identical actions without follow-up, and debugging anti-patterns, such as generating a fix without testing it.
  • Semantic alignment between thoughts and actions is essential: Even rare misalignments strongly correlate with failure or increased computational cost, underscoring the need for explicit validation mechanisms (e.g., self-reflection or critique) to ensure coherence.
@@
    Methodology. We introduce a novel trajectory-based methodology to study the interplay of thoughts, actions, and results in software engineering agents.
@@
    Empirical results. We provide the first systematic study of agent trajectories across multiple state-of-the-art software engineering agents, revealing characteristic patterns and challenges.
@@ II-A Data Collection
We collect trajectory logs from three state-of-the-art software engineering agents:
  • RepairAgent [^1] https://github.com/sola-st/RepairAgent [14]: An autonomous agent for program repair evaluated on 835 bugs of Defects4J [32]. We consider the main trajectories of RepairAgent and we ignore the mutation sub-trajectory
@@
To ensure a diverse yet manageable dataset for detailed manual annotation, we randomly sample 40 trajectories per agent. Given the average success rate of approximately 27% across the three agents, we proportionally sample about 10 successful trajectories per agent. Each trajectory captures a complete bug repair or issue resolution through iterative agent-environment interactions. Our dataset thus comprises 120 trajectories totaling 2,822 iterations across all agents.
@@ II-C Statistical Analysis
The trajectory length may be affected by various factors, including task complexity, effort spent by an agent, and the verbosity of the agent’s reasoning. To manage computational effort, agents limit the trajectory length, e.g., RepairAgent caps at 40 iterations, while OpenHands allows up to 100 iterations.
@@
$\sum_{i=1}^{n}\left(\mathit{tokens}(t_{i})+\mathit{tokens}(a_{i})+\mathit{tokens}(input_{i})\right)$
where $\mathit{tokens}()$ denotes the number of tokens consumed by or generated by the LLM and $\mathit{input_{i}}$ denotes the input prompt at iteration $\mathit{i}$, which integrates $\mathit{r_{i}}$.
@@
We define success as reported in the original evaluation of the agents, i.e., whether the agent produced a valid patch or solution to the task. Specifically, for AutoCodeRover and OpenHands, we consider a trajectory successful if the agent produces a patch that passes the tests executed by SWE-bench [33]. For RepairAgent, we consider a trajectory successful if the produced patch passes the tests executed by Defects4J [32] (called “plausible”) and manual inspection confirms that the patch is equal or semantically to the developer-produced patch (called “correct”).
@@ II-D Categorizing Actions
To categorize the actions, we first map known tools used by agents, such as search tools into the respective category (e.g., Search). Then, we manually inspect the remaining actions that we were unable to map automatically (e.g., a terminal command). The labeling is initially performed by one author, followed by a discussion of ambiguous and uncertain cases to resolve them collaboratively among both authors. A small subset of actions (8.3%) does not fit into any of the above categories, such as internal state transitions in RepairAgent and LLM-suggested actions that AutoCodeRover fails to parse. We exclude these cases from subsequent analyses.
@@ II-E Mining Sequential Action Patterns
We analyze sequences of categorized actions in agent trajectories to identify recurrent decision-making patterns. Each trajectory is represented as a sequence of action labels, e.g., Explore $\rightarrow$ Locate $\rightarrow$ Generate Fix $\rightarrow$ Reproduce $\rightarrow$ Run Tests. We study these sequences by mining for frequently occurring subsequences. First, we implement a lightweight Python script to extract fixed-length $n$-grams and compute their frequencies, choosing $n$ large enough to avoid trivial short sequences, but small enough to retain sufficient diversity and easy-to-analyze patterns. From our experiments, $n\in\{4,5,6\}$ performs well, and for this study we set $n=4$. Next, for each agent we identify the most frequent 4-grams and analyze how their relative frequencies differ between successful and failed tasks. Finally, we flag patterns overrepresented in failing trajectories, as well as rare or unusual 4-grams that may indicate nonstandard or suboptimal workflows.
@@ II-F Labeling Semantic Relationships
Thought $t_{i}$ $\rightarrow$ action $t_{i}$ (within one iteration):
@@
Thought $t_{i}$ $\rightarrow$ thought $t_{i+1}$ (across consecutive iterations):
@@
Action $a_{i}$ $\rightarrow$ action $a_{i+1}$ (across consecutive iterations):
@@
Result $r_{i}$ $\rightarrow$ thought $t_{i+1}$ (across consecutive iterations):
@@
Result $r_{i}$ $\rightarrow$ action $a_{i+1}$ (across consecutive iterations):
@@
These relationships capture how the agent’s results influence its subsequent actions. For example, we label a pair of result and action as Informative when the result provides information that is incorporated into the next action (e.g., an API name found through search).
@@
The relationships in Table I are the result of labeling 40 trajectories from each of the three agents, totaling 14K pairs of trajectory components. The annotation process was initially conducted by the first author over five months of full-time work. To ensure quality and consistency, both authors collaborated closely to discuss the set of labels, resolve ambiguous and difficult cases, and cross-validate individual annotations. Moreover, the main annotator periodically revisited previously labeled examples to ensure consistency across trajectories and re-investigated abnormal numbers (e.g., high misalignment labels in an agent).
@@ III-A RQ1: Trajectory-Level Properties
We investigate trajectory properties aiming to understand how their iterative processes differ in length and complexity. We answer this question by quantitatively analyzing the distribution of iteration counts and token consumption.
Figure 2: Comparative trajectory length between successful and unsuccessful trajectories. Note that the y-axis differs between the agents.
@@
The figure also compares trajectory lengths between successful and unsuccessful instances, revealing agent-specific patterns. For RepairAgent, unsuccessful cases have longer trajectories (mean 40) than successful ones (mean 22), indicating higher effort on difficult tasks, consistent with its design to continue attempts until the agent finds a fix or exceeds the budget. OpenHands shows a similar trend, though a bit less pronounced, with some unsuccessful trajectories reaching the 100-iteration cap. For AutoCodeRover, successful and unsuccessful trajectories have similar average lengths, but unsuccessful ones exhibit higher variability and outliers, often due to internal errors like Response Parse Error. AutoCodeRover terminates after patch application without integrating tests into its workflow.
@@
Figure 3: Input, Output, and Total tokens over successful trajectories and unsuccessful ones. The y-axis differs for each agent.
@@
  • AutoCodeRover (ACR): Trajectories are consistently short and efficient. Failures often involve Error Response Parsing sequences that modestly increase token usage.
@@ III-B1 Overall Action Usage
Figure 4(a) compares the overall distribution of action categories. Across all agents, the most frequent actions are Generate Fix (23%), Run tests (19%), Search (15%), and Explore (14%). While these core actions are common, each agent exhibits distinct patterns in action selection and sequencing.
RepairAgent tends to emphasize a balanced approach incorporating search and exploration, fix generation, and comprehensive testing. In contrast, OpenHands places particular emphasis on a test-driven methodology, including the generation of issue-reproducing tests, which is distinctive to OpenHands’s workflow. AutoCodeRover operates through a structured search–locate–fix process, typically omitting explicit Reproduce and Test steps in its main pipeline. It is worth noting that a subsequent extension, SpecRover [36], introduces dedicated reproduction and testing phases.
@@ III-B2 Action Usage Over Task Progress
To analyze behavior evolution, we normalize iterations $i\in[1..N]$ of a trajectory to percentage progress. Figures 4(b)4(d) depict the proportion of actions per category at each progress point (e.g., 20% of the total trajectory length). In RepairAgent (Fig. 4(b)), the initial 0–20% of the trajectory is evenly split between Fix Generation plus Running Tests ( 50%) and the remaining categories (Explore, Explain, Search, Locate). From 20% to 100%, Fix Generation and Running Tests dominate ( 70%). OpenHands (Fig. 4(c)) starts with  50% Explore and Search in the first 20%, then shifts towards increased Fix Generation, Reproduce, and Running Tests, which constitute  75% in the last 20%. AutoCodeRover (Fig. 4(d)), with median $N=5$, spends the first 50% on Search, Locate, and Explain, then transitions to Locate (30%) and Generate Fix (20%) in the latter half. The sharp transitions in AutoCodeRover’s action distribution are due to its short trajectories. Failing trajectories include many ErrorResponse events, omitted from the plot.
@@ III-B3 Sequence Patterns
To complement the time-distributed action plots, Figures 5, 6, 7 presents a visualization of the most frequent 4-gram sequences executed by each agent. Each row represents a distinct sequence and its corresponding frequency in successful and unsuccessful trajectories.
In the case of RepairAgent, successful 4-grams prominently feature cycles that alternate between Generate Fix and exploratory actions, such as Explain or Explore, with little dominance by any single pattern, as shown in Figure 5. In contrast, unsuccessful trajectories often show many repetitive Generate FixRunning Tests cycles, as shown in the last row of Figure 5.
For OpenHands (Figure 6), the most frequent successful sequences balances context gathering, patch generating, and testing (both generating and running tests). Unsuccessful trajectories, on the other hand, show more repetitive cycles of exploration and search actions (e.g., first rows in the figure), or sequences of repeated testing and fixing steps without interleaving actions. This highlights the importance balancing different actions and enforcing different debugging phases instead of continuously repeating the same actions.
@@
Figure 5: Top 11 sequences of action 4-grams (RepairAgent).
Figure 6: Top 12 sequences of action 4-grams (OpenHands).
Figure 7: Top 11 sequences of action 4-grams (AutoCodeRover).
@@
1) Repeated identical actions without follow-up: For instance, repeated search actions occur across all three agents. While some repetitions arise from unsuccessful prior searches, others neglect to act on the found results. RQ3 elaborates on this no-influence phenomenon.
2) Repeated fix generation without testing: As observed mainly in OpenHands and AutoCodeRover, agents sometimes generate fixes without testing them. RepairAgent mandates testing after fix generation, as it is crucial to validate that a fix works and identify any necessary patch modifications.